### v2

# Tightening constraints on primordial oscillations with latest ACT and SPT data

Abstract
The oscillation feature in primordial power spectrum (PPS), a fingerprint of not only a wide class of models of inflation but new physics, is of significant theoretical interest, and can be imprinted on the cosmic microwave background (CMB). In this work, we present constraints on periodic oscillations in the PPS using the latest ACT DR6 and SPT-3G D1 CMB data with the precise measurements at high multipoles beyond the Planck angular resolution and sensitivity. It is found that the combination of SPT and ACT with Planck CMB dataset significantly tightens the upper bound to $A_{\mathrm{log,lin}}\lesssim 0.029$ at $95\%$ C.L., showing no hint for primordial oscillations, where $A_{\mathrm{log,lin}}$ are the amplitudes of logarithmic and linear oscillation in the PPS, respectively. Our work presents state-of-the-art CMB constraints on primordial oscillations, highlighting the power of the ground-based CMB experiments in constraining physics beyond the simplest slow-roll models.
Introduction.— Inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982); Starobinsky (1980), the standard paradigm of the very early universe, predicts a nearly scale-invariant primordial power spectrum (PPS) of scalar perturbations. Recent Planck cosmic microwave background (CMB) data Aghanim et al. (2020a) strongly support the simplest slow-roll inflationary scenarios, and its combination with recent BICEP/Keck data Ade et al. (2021) has placed stringent constraints on the landscape of models Akrami et al. (2020).
However, physics beyond simplest slow-roll potentials Starobinsky (1992); Adams et al. (2001); Wang et al. (2005); Gong (2005); Slosar et al. (2019); Achúcarro et al. (2022) can result in the oscillations in the PPS, which also are consistent with current data, see also models with null energy condition violation Cai and Piao (2021, 2022); Cai et al. (2024). In particular, the periodic oscillation feature is of significant theoretical interest, motivated by such as Trans-Planckian effects Martin and Brandenberger (2001); Easther et al. (2002); Danielsson (2002); Baumann and McAllister (2015), axion monodromy inflation in string theory Silverstein and Westphal (2008); McAllister et al. (2010); Flauger et al. (2010), and primordial standard clocks Chen (2012); Chen et al. (2016, 2019). There have been some earlier searches for such oscillations with CMB data Easther et al. (2005); Meerburg et al. (2014); Meerburg and Spergel (2014); Peiris et al. (2013); Easther and Flauger (2014); Aich et al. (2013). Currently, the most stringent constraints on primordial oscillations were reported by the Planck cooperation based on their CMB data Akrami et al. (2020), $A_{\mathrm{log,lin}}\lesssim 0.038$ at $95\%$ C.L., where $A_{\mathrm{log,lin}}$ are the amplitudes of logarithmic and linear oscillation in the PPS respectively, which recently has been complemented by constraints from large-scale structure (LSS) surveys, e.g.Beutler et al. (2019); Ballardini et al. (2023); Mergulhão et al. (2023). See also Braglia et al. (2021, 2022a, 2022b, 2023); Petretti et al. (2025) for constraints on other realistic models of primordial features.
[…]
In this work, we present state-of-the-art constraints on primordial oscillations leveraging the latest ACT DR6 and SPT-3G D1 data. As shown in Fig. 1, we find that the constraints set by SPT+ACT on primordial oscillations are comparable to those from Planck, and the combined Planck+SPT+ACT dataset significantly tightens the upper bound to $A_{\mathrm{log,lin}}\lesssim 0.029$ at $95\%$ C.L., showing no hint for primordial oscillations.
Figure 1: 1D posterior distributions of the logarithmic (left) and linear (right) oscillating amplitudes, obtained using Planck, SPT+ACT, and Planck+SPT+ACT datasets, respectively.
Primordial oscillations.— Typically, different new physics beyond the simplest slow-roll models can make the PPS manifest different oscillatory patterns. The sudden changes in the inflaton potential, such as bumps and steps Starobinsky (1992); Adams et al. (2001); Gong (2005); Hazra et al. (2010) or sharp turns Ashoorioon et al. (2009); Achucarro et al. (2011); Konieczka et al. (2014), can imprint linear oscillations onto the PPS, while the resonant models Chen et al. (2008) and the axion monodromy model Silverstein and Westphal (2008); McAllister et al. (2010); Flauger et al. (2010) can lead to logarithmically oscillating signals in the PPS.
The templates we consider for the corresponding linear and logarithmical oscillating PPS are, e.g.Akrami et al. (2020); Slosar et al. (2019), as shown in Fig. 2:
$P_{\mathcal{R}}(k)=P_{\mathcal{R},0}(k)\left[1+A_{l}\cos\left(\omega_{l}({\frac{k}{k_{*}}})_{l}+\phi_{l}\right)\right],$ (1)
where $P_{\mathcal{R},0}(k)=A_{\mathrm{s}}(k/k_{*})^{n_{\mathrm{s}}-1}$ is the standard power-law PPS, the pivot scale is $k_{*}=0.05\,\mathrm{Mpc}^{-1}$, and $l=\mathrm{log}$ and $({k/k_{*}})_{\mathrm{log}}=\log(k/k_{*})$ for logarithmic oscillation, while $l=\mathrm{lin}$ and $({k/k_{*}})_{\mathrm{lin}}=k/k_{*}$ for linear oscillation.
Figure 2: Primordial power spectra with the logarithmic and linear oscillations, where $A_{lin,log}=0.05$, $\omega_{lin,log}=10$, and $\phi_{lin,log}=3.14$.
Datasets and Methods.— In our analysis, inspired by Camphuis et al. (2025), we consider three datasets as follows:
  • Planck: The Plik-lite likelihood for Planck PR3 high-$\ell$ temperature and polarization data, and the Commander likelihood for low-$\ell$ TT and the SimALL likelihood for low-$\ell$ $E$ Aghanim et al. (2020b), as well as the Planck PR4 CMB lensing data Carron et al. (2022).
  • SPT+ACT: The foreground-marginalized SPT-lite [^1] https://github.com/SouthPoleTelescope/spt_candl_data likelihood for SPT-3G D1 Camphuis et al. (2025); Balkenhol et al. (2024), combined with the ACT-lite [^2] https://github.com/ACTCollaboration/DR6-ACT-lite likelihood for ACT DR6 Louis et al. (2025); Calabrese et al. (2025), as well as the ACT DR6 CMB lensing data [^3] We use the actspt3g_baseline variant of https://github.com/qujia7/spt_act_likelihood. Madhavacheril et al. (2024); Qu et al. (2024); MacCrann et al. (2024) and SPT-3G Ge et al. (2025); Qu et al. (2025). Here, we impose a Gaussian prior on the optical depth, $\tau_{\mathrm{reio}}=0.054\pm 0.007$ Aghanim et al. (2020a).
  • Planck+SPT+ACT: The combination of Planck and SPT+ACT, as well as the CMB lensing data [^4] We use the actplanckspt3g_baseline variant of https://github.com/qujia7/spt_act_likelihood. Following Refs.Louis et al. (2025); Camphuis et al. (2025), we cut the Planck CMB data to $\ell_{max}=1000$ for TT and $\ell_{max}=600$ for TE and EE.
In addition, we include the DESI DR2 baryon acoustic oscillations (BAO) dataset Abdul Karim et al. (2025) and the Pantheon+ supernova (SN) dataset Scolnic et al. (2022).
Parameter Prior
$A_{X}$ $[0,0.5]$
$\omega_{X}$ $[0,100]$
$\phi_{X}$ $0,2\pi]$
$\log(10^{10}A_{\mathrm{s}})$ $[1.61,3.91]$
$n_{\mathrm{s}}$ $[0.8,1.2]$
$H_{0}$ $[20,100]$
$\Omega_{\mathrm{b}}h^{2}$ $[0.005,0.1]$
$\Omega_{\mathrm{c}}h^{2}$ $[0.001,0.99]$
$\tau_{\mathrm{reio}}$ $[0.01,0.8]$
Table 1: The priors for relevant parameters in our MCMC analysis. Note that a Gaussian prior, $\tau_{\mathrm{reio}}=0.054\pm 0.007$, is applied only to the SPT+ACT dataset.
We perform the Markov chain Monte Carlo (MCMC) analysis using Cobaya Torrado and Lewis (2021). The observables are computed using the cosmological Boltzmann code CLASS Blas et al. (2011). The Gelman-Rubin criterion Gelman and Rubin (1992) for all chains is converged to R1<0.02R-1<0.02. The flat priors adopted for relevant parameters are presented in Table 1. The best-fit parameters and corresponding $\chi^{2}$ values are obtained using the global optimization method in PROSPECT Holm et al. (2023).
Results.— In Table 2, we see that compared to the Planck results, SPT+ACT yields a slightly tighter constraint on the logarithmic oscillating amplitude $A_{\mathrm{log}}$, but a higher upper limit on $A_{\mathrm{lin}}$ due to the long posterior tail. The combined Planck+SPT+ACT dataset yields our most stringent constraints, $A_{\mathrm{log}}<0.0286$ and $A_{\mathrm{lin}}<0.0267$ ($95\%$ C.L.).
Dataset Planck SPT+ACT Planck+SPT+ACT
$A_{\mathrm{log}}$ <0.0387<0.0387 <0.0335<0.0335 <0.0286<0.0286
$A_{\mathrm{lin}}$ <0.0357<0.0357 <0.0433<0.0433 <0.0267<0.0267
Table 2: The $95\%$ upper limits on the oscillating amplitudes from different datasets.
Figure 3: The 2D posterior distribution of $A_{l}$-$\omega_{l}$ ($68\%$ and $95\%$ confidence levels) for logarithmic (upper) and linear (lower) oscillating spectra, obtained using Planck, SPT+ACT, and Planck+SPT+ACT datasets, respectively.
In Fig. 3, we show the posterior distributions of $A_{\mathrm{log,lin}}$-$\omega_{\mathrm{log,lin}}$. Some localized $A_{\mathrm{log,lin}}$ peaks can be observed at specific frequencies, but they do not match between the results of Planck and SPT+ACT. This suggests that these apparent peaks are more likely statistical fluctuations rather than real physical signals. The combined Planck+SPT+ACT dataset tightens the constraints across all frequencies, with the $2\sigma$ upper limits consistently below $0.05$.
Dataset Planck SPT+ACT
Planck+
SPT+ACT
log lin log lin log lin
low-$\ell$ $T$ 0.43-0.43 0.33-0.33 $0.07$ 0.48-0.48
low-$\ell$ $E$ $0.41$ $0.00$ 0.27-0.27 1.32-1.32
Planck high-$\ell$ 8.04-8.04 10.33-10.33 2.74-2.74 2.20-2.20
SPT-3G D1 3.73-3.73 $0.08$ 0.57-0.57 1.25-1.25
ACT DR6 2.68-2.68 6.79-6.79 4.84-4.84 1.24-1.24
CMB lensing $0.51$ 0.00-0.00 0.19-0.19 0.11-0.11 $0.04$ 0.39-0.39
DESI BAO 1.08-1.08 $0.33$ 0.32-0.32 $0.26$ $0.13$ $0.56$
Pantheon+ SN $0.38$ $0.09$ 0.04-0.04 $0.06$ 0.01-0.01 0.11-0.11
Total $\Delta\chi^{2}$ $\mathbf{-8.26}$ $\mathbf{-10.97}$ $\mathbf{-6.11}$ $\mathbf{-7.16}$ $\mathbf{-8.33}$ $\mathbf{-6.43}$
Preference level $\mathbf{1.7\sigma}$ $\mathbf{2.3\sigma}$ $\mathbf{1.3\sigma}$ $\mathbf{1.5\sigma}$ $\mathbf{1.8\sigma}$ $\mathbf{1.3\sigma}$
Table 3: $\Delta\chi^{2}$ values for the best-fit models with logarithmic and linear oscillations, relative to models with the standard PPS, obtained using the Planck, SPT+ACT, and Planck+SPT+ACT datasets, respectively. Note that the CMB lensing data used for each dataset is different.
To evaluate the statistical significance for primordial oscillations, we calculate the $\Delta\chi^{2}$ values for the best-fit models of the oscillating PPS relative to the standard PPS, as summarized in Table 3. The Planck data indicate some hints for primordial oscillations, we have $\Delta\chi^{2}_{\mathrm{log}}=-8.26$ and $\Delta\chi^{2}_{\mathrm{lin}}=-10.97$ for logarithmic and linear oscillations, respectively. This corresponds that primordial oscillations are preferred at the $1.7\sigma$ and $2.3\sigma$ significance levels. However, for SPT+ACT, the corresponding values are suppressed to $\Delta\chi^{2}_{\mathrm{log}}=-6.11$ ($1.3\sigma$) and $\Delta\chi^{2}_{\mathrm{lin}}=-7.16$ ($1.5\sigma$). The further inclusion of Planck data into SPT+ACT does not lead to any significant $\Delta\chi^{2}$ shift, $\Delta\chi^{2}_{\mathrm{log}}=-8.33$ ($1.8\sigma$) and $\Delta\chi^{2}_{\mathrm{lin}}=-6.43$ ($1.3\sigma$). Therefore, essentially latest SPT+ACT dataset suppress the oscillations in the PPS hinted by Planck, so no hints for primordial oscillations from the Planck+SPT+ACT datasets.
Figure 4: The CMB spectral residuals of the best-fit models with logarithmic and linear oscillations obtained using Planck (dashed) and Planck+SPT+ACT (solid), respectively. The base line is the Planck best-fit model without oscillations. The ACT and SPT-3G data points are also showed.
In Fig. 4, we show the spectral residuals of the best-fit models with oscillations in the PPS for Planck (dashed) and Planck+SPT+ACT (solid) datasets respectively, relative to the Planck best-fit model with standard power-law spectrum. As expected, the impact of the primordial oscillations is prominent in the small-scale polarization spectra, and there is a slight suppression of the residuals around $l\sim 500-1000$ in the TE and EE spectra for Planck+SPT+ACT compared to Planck results.
Conclusion.— The search for primordial oscillations is of significant theoretical interest and would open a crucial observational window into the physics of inflation beyond the simplest slow-roll models. In this work, we have reported the constraints on primordial oscillations with the latest ACT DR6 and SPT-3G D1 datasets. The combination of SPT and ACT with Planck dataset yield our most stringent constraints, the upper limits on the oscillating amplitudes of $A_{\mathrm{log}}<0.0286$ and $A_{\mathrm{lin}}<0.0267$ ($95\%$ C.L.).
Our work presents state-of-the-art CMB constraints on primordial oscillations, highlighting the power of the ground-based CMB experiments, especially when combined with large-scale Planck data. It is expected that upcoming CMB observations, such as the Simons Observatory Ade et al. (2019) and CMB-S4 Abazajian et al. (2019) will further improve our constraints on primordial oscillations and bring deeper insights into the physics of inflation. It is well known that LSS also offer a vital, independent probe to primordial oscillations, especially those from the full-shape galaxy clustering measurements, e.g. see Refs. Ballardini et al. (2023); Mergulhão et al. (2023); Calderon et al. (2025), thus it will be promising to revisit the constraints on primordial oscillations by using a combination of the corresponding DESI Adame et al. (2024) and Euclid Laureijs et al. (2011) data with CMB dataset.
[…]
Acknowledgements.
This work is supported by NSFC, No.12475064, 12075246, National Key Research and Development Program of China, No. 2021YFC2203004, and the Fundamental Research Funds for the Central Universities. We acknowledge the use of high performance computing services provided by the International Centre for Theoretical Physics Asia-Pacific cluster and the Tianhe-2 supercomputer.
Supplemental Material for “Tightening constraints on primordial oscillations with latest ACT and SPT data”
I Full results of MCMC and best-fit parameters
The full results of our MCMC analysis, including the best-fit parameters and $\chi^{2}$ values for models with primordial oscillations, are presented in the Tables 1, 2 and 3. We can see that the inclusion of primordial oscillations has negligible impact on the results of other cosmological parameters.
Parameter Planck SPT+ACT Planck+SPT+ACT
$10^{9}A_{\mathrm{s}}$ $2.115(2.113)\pm 0.029$ $2.142(2.163)\pm 0.024$ $2.141(2.142)^{+0.025}_{-0.029}$
$n_{\mathrm{s}}$ $0.9702(0.9690)\pm 0.0033$ $0.9739(0.9751)\pm 0.0057$ $0.9734(0.9732)\pm 0.0030$
$H_{0}$ $68.31(68.25)\pm 0.29$ $68.09(68.19)\pm 0.26$ $68.11(68.14)\pm 0.25$
$100\Omega_{\mathrm{b}}h^{2}$ $251(2.248)\pm 0.012 2.2$$47(2.247)\pm 0.013 2.2$$2455(2.2455)\pm 0.0090$
$\Omega_{\mathrm{c}}h^{2}$ $0.11783(0.011792)\pm 0.00064$ $0.11819(0.11785)\pm 0.00066$ $0.11801(0.11792)\pm 0.00061$
$\tau_{\mathrm{reio}}$ $0601(0.0558)\pm 0.0070 0.0$$646(0.0691)\pm 0.0059 0.0$$0650(0.0641)^{+0.0066}_{-0.0075}$
$\chi^{2}$ $2231.40$ $1775.00$ $2420.51$
Table 1: The mean (best-fit) $\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with the standard PPS.
Parameter Planck SPT+ACT Planck+SPT+ACT
$A_{\mathrm{log}}$ <0.0387(0.0196)<0.0387(0.0196) <0.0335(0.0153)<0.0335(0.0153) <0.0286(0.0145)<0.0286(0.0145)
$\omega_{\mathrm{log}}$ unconstrained(31.46)(31.46) unconstrained(36.23)(36.23) unconstrained $(31.44)$
$\phi_{\mathrm{log}}$ unconstrained(4.39)(4.39) unconstrained(6.13)(6.13) unconstrained(4.23)(4.23)
$10^{9}A_{\mathrm{s}}$ $2.113(2.114)\pm 0.029$ $2.144(2.162)\pm 0.025$ $2.141(2.140)\pm 0.027$
$n_{\mathrm{s}}$ $0.9703(0.9716)\pm 0.0033$ $0.9733(0.9728)\pm 0.0060$ $0.9733(0.9727)\pm 0.0029$
$H_{0}$ $68.31(68.44)\pm 0.29$ $68.11(68.24)\pm 0.26$ $68.11(68.16)\pm 0.25$
$100\Omega_{\mathrm{b}}h^{2}$ $248(2.259)\pm 0.013 2.2$$49(0.2257)\pm 0.013 2.2$$2462(2.2498)\pm 0.0097$
$\Omega_{\mathrm{c}}h^{2}$ $0.11783(0.11766)\pm 0.00063$ $0.11817(0.11805)\pm 0.00065$ $0.11803(0.11798)\pm 0.00061$
$\tau_{\mathrm{reio}}$ $0597(0.0579)\pm 0.0069 0.0$$646(0.0686)\pm 0.0060 0.0$$0648(0.634)\pm 0.0070$
$\chi^{2}$ $2423.140$ $1768.89$ $2412.18$
Table 2: The mean (best-fit) $\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with logarithmic oscillations. For upper limits, we report the one-side $95\%$ confidence level ($2\sigma$).
Parameter Planck SPT+ACT Planck+SPT+ACT
$A_{\mathrm{lin}}$ <0.0357(0.0136)<0.0357(0.0136) <0.0433(0.0086)<0.0433(0.0086) <0.0267(0.0229)<0.0267(0.0229)
$\omega_{\mathrm{lin}}$ unconstrained(10.92)(10.92) $56(5.41)^{+40}_{-50}$ unconstrained(75.23)(75.23)
$\phi_{\mathrm{lin}}$ unconstrained(4.01)(4.01) unconstrained(6.20)(6.20) unconstrained(6.28)(6.28)
$10^{9}A_{\mathrm{s}}$ $2.113(2.104)^{+0.027}_{-0.031}$ $2.142(2.159)\pm 0.024$ $2.141(2.126)^{+0.024}_{-0.028}$
$n_{\mathrm{s}}$ $0.9701(0.9706)\pm 0.0033$ $0.9742(0.9766)\pm 0.0057$ $0.9734(0.9728)\pm 0.0030$
$H_{0}$ $68.30(68.26)\pm 0.29$ $68.10(68.22)\pm 0.26$ $68.10(68.11)\pm 0.25$
$100\Omega_{\mathrm{b}}h^{2}$ $250(2.243)\pm 0.013 2.2$$47(2.247)\pm 0.012 2.2$$2457(2.246)\pm 0.0094$
$\Omega_{\mathrm{c}}h^{2}$ $0.11784(0.11780)\pm 0.00063$ $0.11816(0.11785)\pm 0.00067$ $0.11803(0.11805)\pm 0.00061$
$\tau_{\mathrm{reio}}$ $0594(0.0559)^{+0.0064}_{-0.0074} 0.0$$645(0.0693)\pm 0.0059 0.0$$0649(0.0605)^{+0.0062}_{-0.0073}$
$\chi^{2}$ $2420.42$ $1767.84$ $2414.09$
Table 3: The mean (best-fit) $\pm 1\sigma$ errors of relevant parameters and $\chi^{2}$ values for models with linear oscillations. For upper limits, we report the one-side $95\%$ confidence level ($2\sigma$).